What is the volume of the largest ball that will fit inside a cubical box that has a volume of 64 cubic inches?
Answer by stanbon(48526)

What is the volume of the largest ball that will fit inside a cubical box that has a volume of 64 cubic inches?
----------------
Volume of the cube = side^3 = 64
side = cbrt(64) = 4
----------------------
And diameter of the ball = the length of a side = 4 inches
Volume of the ball = (4/3)pir^3 = (4/3)pi(2^3) = (32/3)pi in^3
==================================================================
Cheers,
Stan H.What is the volume of the largest ball that will fit inside a cubical box that has a volume of 64 cubic inches?
===============================
Cheers,
Stan H.